the sum of 4 times a number and -2 is equal to the sum of 5 times the number and -2
Answer by oberobic(2304)

The answer depends on how you interpret the statements.
.
It could be:
.
4x +(-2) = 5x -2
4x -2 = 5x -2
4x = 5x
which is only true when x=0
.
But the equation could be:
.
4x +(-2) = 5(x -2)
4x -2 = 5x -10
-x = -8
x = 8
.
check the answer
.
4(8) -2 = 30
5(8-2) = 30
Correct.
.
Done.